OpenStudy (anonymous):

how do you find the limit of x^sinx as x approaches 0 from the right?

OpenStudy (anonymous):

z = lim (sin x)^x . . .x → 0 ln z = lim x ln (sin x) . . .x → 0 ln z = lim ln (sin x)/(1/x) . . .x → 0 using L'Hospital's rule, ln z = - lim (cos x)(1/sin x)/(1/x²) . . .x → 0 ln z = - lim x (cos x)(x/sin x) . . .x → 0 ln z = (0) (1)(1) = 0 z = e^(0) z = 1

OpenStudy (kc_kennylau):

it's \(x^{\sin{x}}\) not \(\sin^xx\)